OpenStudy (anonymous):

f(x)=(7x)/((1+5x^2)^2) ,at an interval -5 is less then or equal to x, and x is less than or equal to 2. The absolute maximum value is _____ and this occurs at x equals _____. The absolute minimum value is _____ and this occurs at x equals _____.

OpenStudy (zzr0ck3r):

you took the derivative?

OpenStudy (anonymous):

I honestly don know how to take the derivative of the function. \[f(x)=\frac{ 7x }{ (1+5x^2)^2 }\]

OpenStudy (zzr0ck3r):

what is the derivative of the top?

OpenStudy (anonymous):

7?

OpenStudy (zzr0ck3r):

bottom?

OpenStudy (anonymous):

2(1+5x^2)*(10x)?

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

yep

OpenStudy (zzr0ck3r):

ok

OpenStudy (zzr0ck3r):

derivative of top *bottom - derivative of bottom * top all over the bottom squared

OpenStudy (zzr0ck3r):

sorry on tablet this is hard

OpenStudy (zzr0ck3r):

\[\frac{7 *((1+5x^2)^2)-7x*2(1+5x^2)*(10x)?}{(((1+5x^2)^2)^2}\]

OpenStudy (zzr0ck3r):

disregard the ?

OpenStudy (anonymous):

\[\frac { (7)((1+5x^2)^2)-(2(1+5^2)^2)(10x)(7x) } { (1+5x^2)^2)^2 }\]

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

oh yeah that's what I got hah

OpenStudy (zzr0ck3r):

ok now we are setting this thing = 0 so we can disregard the bottom

OpenStudy (zzr0ck3r):

so simplify the top

OpenStudy (zzr0ck3r):

wait wait

OpenStudy (zzr0ck3r):

there is a problem with the bottom derivative

OpenStudy (zzr0ck3r):

((1+5x^2)^2 so 2(1+5x^2)*10x

OpenStudy (zzr0ck3r):

we loose the ^2 on the outside

OpenStudy (anonymous):

ohh. okay.

OpenStudy (zzr0ck3r):

ok so now what do you have on top?

OpenStudy (anonymous):

\[(7)((1+5x^2)^2)-(2(1+5x^2))(10x)(7x)\]

OpenStudy (zzr0ck3r):

bah im sorry man its so hard to see on this thing your derivative is fine

OpenStudy (zzr0ck3r):

...

OpenStudy (zzr0ck3r):

ok the problem is we need to set this thing to = 0 and this is a 4th degree polynomial

OpenStudy (zzr0ck3r):

7∗((1+5x 2 )^2−7x∗2(1+5x 2 )∗(10x) = 0 is hard to solve

OpenStudy (anonymous):

I was just going to say the same.

OpenStudy (zzr0ck3r):

factor 1+5x^2 out (1+5x^2)(7(1+5x^2)-140x^2)=0

OpenStudy (zzr0ck3r):

solve 5x^2+1=0 and this solution is complex so we disregard it

OpenStudy (zzr0ck3r):

these solutions are complex*...

OpenStudy (anonymous):

So what do we do now?

OpenStudy (zzr0ck3r):

(7(1+5x^2)-140x^2) = 0 7+35x^2-140x^2 = 0 7-105x^2=0 x= +- sqrt(7/105)

OpenStudy (anonymous):

so how do i find the max and min.and when it occurs as x?

OpenStudy (zzr0ck3r):

one of those is the max and one is the min so take third derivative to find concavity (this is not going to be fun) or you can plug in numbers right next to those to tell sqrt(7/105) is about .25 so plug in .24, sqrt(7/105) and .26 you will get some value for sqrt(7/105) and if the two other numbers are lower, then sqrt(7/105) is a max, and vice versa

OpenStudy (zzr0ck3r):

or do you want to take the second derivative?

OpenStudy (zzr0ck3r):

that should say second derivative two comments ago (not third)

OpenStudy (zzr0ck3r):

you could also graph it and look at it:) http://www.wolframalpha.com/input/?i=graph+%287x%29%2F%28%281%2B5x%5E2%29%5E2%29+

OpenStudy (zzr0ck3r):

how do you want to do this @pdd21

OpenStudy (anonymous):

I want to figure out the answer to the question the easiest way possible (: lol @zzr0ck3r

OpenStudy (zzr0ck3r):

well, do you have to show work?

OpenStudy (zzr0ck3r):

if you understand my argument concerning taking the points .24, sqrt(7/105), .26 then use that. its valid because we know that we only have two real places where slope = 0

OpenStudy (anonymous):

no I dont it's online hw. I plug in an answer and it tells me if i'm from or right. so far I've tried this problem 12 times and still none correct -.-

OpenStudy (zzr0ck3r):

what do you get when you evaluate the original function at sqrt(7/105)?

OpenStudy (anonymous):

I got the max! (: just the min would it be the same values but neg?

OpenStudy (zzr0ck3r):

no

OpenStudy (zzr0ck3r):

sec

OpenStudy (zzr0ck3r):

ps for the future http://www.wolframalpha.com/input/?i=extrema+%287*x%29%2F%281%2B5x%5E2%29%5E2

OpenStudy (anonymous):

ohhhh okay. lol and thanksss so much! (: I really appreciate you helping me! (: @zzr0ck3r

OpenStudy (zzr0ck3r):

np, it looks like something happened in our setting = 0, but im sure you can figure it out. our x is off a little...but it gives the answer.